OpenStudy (anonymous):

sqrt (x+1) +3=sqrt 3x+4 The +3 is on the outside of the sqrt.

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

move the 3 to the right first

OpenStudy (anonymous):

\[\sqrt{x + 1} + 3 = \sqrt{3x + 4} \]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

is the3x + 4 all under the root?

OpenStudy (agreene):

x=15

OpenStudy (anonymous):

I think you should square on both sides.

OpenStudy (anonymous):

can you explain?

OpenStudy (agreene):

assuming, that pratu's is the correct question... thats how I interpreted it as well.

OpenStudy (anonymous):

please explain.

OpenStudy (agreene):

square both sides, expand left, collect likes, factor, square both sides, expand right, collect to the right, complete the square, solve for factors. Is how I did it at least,

OpenStudy (anonymous):

Pratu is correct. I don't get 15 even though that is the correct answer

OpenStudy (mathmagician):

\[(\sqrt{x+1}+3)^2=(\sqrt{3x+4}^2\] \[x+1+6\sqrt{x+1}+9=3x+4\] \[6\sqrt{x+1}=2x-6\] \[36x+36=4x^2-24x+36\] \[4x^2-60x=0\] \[4x(x-15)=0\] x=15

OpenStudy (anonymous):

thanks.

OpenStudy (agreene):

Thats not quite how I did it, but yeah... lol :D

OpenStudy (anonymous):

Where does the (x+1+6) come from in this part of the problem?[x + 1 +6\sqrt{x}+9\] I understand the 9 but not sure where the 6 is from

OpenStudy (mathmagician):

\[(\sqrt{x+1})^2=x+1\] and\[2*3*\sqrt{x+1}=6\sqrt{x+1}\], because \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (agreene):

\[\sqrt{x+1}+3 = \sqrt{3 x+4}\] \[(\sqrt{x+1}+3)^2 = 3 x+4\] \[x+6 \sqrt{x+1}+10 = 3 x+4\] \[6 \sqrt{x+1} = 2 x-6\] \[\sqrt{x+1} = \frac16 (2 x-6)\] \[x+1 = \frac1{36} (2 x-6)^2\] \[x+1 = \frac{x^2}{9}-\frac{2 x}3+1\] \[\frac{5 x}3-\frac{x^2}9 = 0\] \[x^2-15 x = 0\] \[x^2-15 x+\frac{225}4 = \frac{225}4\] \[(x-15/2)^2 = \frac{225}4\] \[|x-15/2|= 15/2\] x-15/2 = -15/2 or x-15/2 = 15/2 x = 0 or x-15/2 = 15/2 x = 0 or x = 15 Looking at the case for 0, you'll see it isnt a solution.

OpenStudy (anonymous):

Thanks to everyone!